Weather:

            Light winds during the night. A few clouds this morning with calm conditions. High humdity

Status:

            TRH at 2m on A17 went down last night. Used TIO command to get it going again.

            The QCtable for Lambdasoil shows bad data however QCtable A11rs shows good data. A data_dump confirms the data is coming in. Killed lambdasoil QC and started another one. Data coming in.  

To Do:

          Finish potting TRH units. Shoot boom angles and clean Kryptons.